OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6.

OpenStudy (anonymous):

help please

OpenStudy (callisto):

What is the slope of the line y = 3x + 6?

OpenStudy (anonymous):

Thats what i dont understand it just has these answers y = -1/3x + 4 y = -3x - 16 y = 3x + 20 y = 1/3x

OpenStudy (callisto):

For an equation of this form y=mx+c m=slope ; c= y-intercept. So, can you identify the slope of the line in that equation now?

OpenStudy (anonymous):

1/3x ?

OpenStudy (callisto):

Why ._.?

OpenStudy (anonymous):

.-. I HAVE 1 minute to answer this

OpenStudy (callisto):

Is it a quiz?

OpenStudy (anonymous):

yes but i didnt really get this in class

OpenStudy (callisto):

1. Identify the slope in the given equation by observation.. (Said above) 2. Slope of the equation = -1/slope of line given (perpendicular) 3. Use point-slope form to write the equation. Please do not use ask quiz question here - Don't use OpenStudy to cheat..

OpenStudy (anonymous):

Okay i apologize. Thank you for your help.

OpenStudy (callisto):

Good luck with your quiz!